# Learning Confident Classifiers in the Presence of Label Noise

### v2

@@ 5 Experiments
5.1 Classification
Noise rate Ours-Inf Ours-Ent Co-tea Co-tea+ JoCoR Trace CDR
@@ 5.1 Classification
Table 1: Comparison of test accuracy (%) (mean $\pm$ st. dev.) for CIFAR-10. Best results are in bold, second-best underlined.
Noise rate Ours-Inf Ours-Ent Co-tea Co-tea+ JoCoR Trace CDR
symmetric 20% 90.67 90.79 90.48 88.69 91.88 90.61 88.69
$\pm$ 0.21 $\pm$ 0.19 $\pm$ 0.06 $\pm$ 0.09 $\pm$ 0.06 $\pm$ 0.24 $\pm$ 0.09
symmetric 30% 91.35 90.34 90.36 88.50 91.33 89.64 87.38
$\pm$ 0.25 $\pm$ 0.26 $\pm$ 0.13 $\pm$ 0.09 $\pm$ 0.08 $\pm$ 0.30 $\pm$ 0.09
symmetric 50% 89.51 89.49 89.37 77.96 89.21 88.94 85.36
$\pm$ 0.55 $\pm$ 0.58 $\pm$ 0.08 $\pm$ 0.33 $\pm$ 0.06 $\pm$ 0.53 $\pm$ 0.09
pairflip 20% 90.90 90.77 90.68 89.12 91.37 90.40 90.01
$\pm$ 0.33 $\pm$ 0.29 $\pm$ 0.05 $\pm$ 0.08 $\pm$ 0.09 $\pm$ 0.35 $\pm$ 0.11
pairflip 30% 90.38 90.65 90.11 89.06 89.67 90.33 88.78
$\pm$ 0.41 $\pm$ 0.45 $\pm$ 0.08 $\pm$ 0.10 $\pm$ 0.07 $\pm$ 0.46 $\pm$ 0.14
pairflip 45% 89.37 89.02 78.86 52.61 88.10 89.08 64.63
$\pm$ 0.68 $\pm$ 0.71 $\pm$ 0.11 $\pm$ 0.43 $\pm$ 0.39 $\pm$ 0.72 $\pm$ 0.17
Noise type Ours-Ent Co-tea Co-tea+ JoCoR Trace CDR
worst 73.58 67.97 65.45 68.74 70.23 59.87
$\pm$ 0.96 $\pm$ 1.17 $\pm$ 0.25 $\pm$ 0.21 $\pm$ 0.89 $\pm$ 0.58
Table 3: Comparison of test accuracy (%) (mean $\pm$ st. dev.) for CIFAR-10N. Best results are in bold, second-best underlined.
5.1.1 Classification Task Baselines
We compare our algorithm with the following approaches
@@ 5.1.1 Classification Task Baselines
5.1.2 Datasets
In this work, we consider dataset with synthetically generated annotations along with annotations from human experts. For classification problem, we consider the standard benchmark datasets: MNIST [11], Fashion-MNIST [52], CIFAR-10 [25] where noises are generated synthetically and CIFAR-10N dataset (real world human annotations noisy data) [48] to demonstrate the effectiveness of our methodology. The noise types, used in the classification experiments, are as follows.
  • The pairflip noise involves swapping the labels of two adjacent categories/classes based on a preset ratio [30].
  • The symmetric noise retains a portion of the original labels and uniformly reassigns the remainder to all other categories [33].
  • The worst is the human label, which comprise of the highest amount of noise.
Detailed descriptions of datasets and examples of noise transition matrices are provided in Appendix B.2.
5.1.3 Performance Evaluation
[…]
5.1.4 Results
[…]
Annotators Original Thin Thick
Annotator 1 symmetric 80% asymmetric 40% pairflip 95%
Annotator 3 pairflip 60% pairflip with permutation 40% symmetric 80%
Table 4: Annotator information for three different styles (MNIST).
Curated Dataset.
We have assembled the dataset that is based on MNIST, where noise level depends on input image style for various annotators. Three type of image styles were simulated by performing morphological transformations (in particular, thinning and thickening) on the original images, using Morpho-MNIST software [8]. In addition to the noise types described in Section 5.2.3, asymmetric and pairflip with permutation where applied. The type and level of noises applied to original labels are provided in Table 4.
For a dataset consisting of three different types of images (original, thin, and thick) and three different annotators (Table 4), we compare (i) base classifier model without annotators and regularization, (ii) our approach without regularization, and (iii) our approach with information-based regularization. Each annotator NN has similar architecture as in classifier model and takes images as an input.
The result of experiments can be seen in Figures 2. Our approach is more accurate and confident compared to the classifier model. The accuracy of our approach with regularizer is higher and more confident than the model without the regularizer.
(a) Testing accuracy
(b) Testing entropy
Figure 2: Accuracy and Entropy for Curated MNIST testing data on seven different seeds.
The proposed approach is able to learn the annotators’ confusion. Predicted confusion matrices for each annotators and different image types are provided in Figure 3. More results can be found in Appendix C.3.2.
Image Original Ours (m=2, ($\lambda$=0.01) Ours ($\lambda$=0)
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Original and Predicted confusion for different Annotators using different models: our approach with regularizer ($\lambda=0.01$, m=2) and without it ($\lambda=0$). (MNIST).
5.2 Segmentation
5.2.1 Segmentation Task Baselines
We compare our approach with the following methods.
@@ 5.2.1 Segmentation Task Baselines
We refer readers to the Appendix, Section C.2 for some additional experiments where we adapt the regularizer from [29] to segmentation problem and compare it to the approaches above.
5.2.2 “Confident regions” (CR) component in the loss
All annotators’ boundaries Initial (d=0) Dilated (d=3)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Visualization of CR with different dilation radius of uncertainty regions (RIGA Optical Cup). Here, uncertainty regions have natural colors, while CR is whitened.
In this work, we also propose to directly apply the segmentation network $\hat{p}_{\theta}(x_{i})$ for predictions for the pixels with low labeling noise. We construct such confident regions (CR) by a special procedure. First, we find such areas of the annotation masks for which all the annotators assign the same label. Next, we use the heuristic that the pixels near the boundary of this region have a high potential for uncertainty. Because of this, we dilate the uncertainty region (complement of the unanimously segmented region) using the convolution operation with zero bias. The complement of this dilated uncertainty region is now a CR, which we will use in our loss function by adding a special term, see equation (3). The process of selecting the CR is illustrated by Figure 4.
5.2.3 Datasets
For segmentation problem we use MNIST [11] as a representative of dataset with artificial noises, where synthetic noisy annotations are created by applying morphological transformations simulations on assumed ground truth [8]. As an example of real world dataset we utilize retinal fundus images for glaucoma analysis (RIGA) [1] and the lung image database consortium (LIDC) [24] datasets. Detailed descriptions of datasets, examples of input and annotations are provided in Appendix B.1, Figures 6, 7, 8. Detailed descriptions of datasets and examples of samples are provided in Appendix B.2.
5.2.4 Performance Evaluation
We performed our experiments using three datasets with real-world and artificial noisy labels. To evaluate the segmentation model, use the Dice coefficient on validation splits and report its mean and standard deviation computed over the last 10 epochs. Note the Dice coefficients are calculated for each sample individually, and then the average value is taken. Although the Dice coefficient between the prediction $\hat{p}_{\theta}(x)$ and true label ${y}$ is more important, we also consider: Dice coefficient among the noisy annotations $y_{a}$ ($D(y_{a},y_{a})$), Dice coefficient among predictions of annotations ($D(\tilde{y}_{a},\tilde{y}_{a})$), Dice coefficient between the annotations’ predictions and noisy labels ($D(\tilde{y}_{a},y_{a})$), Dice coefficient between noisy labels and true labels ($D(y_{a},y_{gt})$), and Dice coefficient between the annotations’ predictions and true label ($D(\tilde{y}_{a},y_{gt})$). More details are provided in Appendix C.1. In the case of medical images, to make the evaluation, we compute the true label by majority voting. We emphasize that they are used only in the evaluation stage, not in the training. Along with these baseline models, we also run the model without involving the confidence region in our framework to see the effect of including it.
5.2.5 Segmentation Results
Datasets MR-Net CM-Net PU-NET Ours (no conf) Ours
MNIST 95.81 93.85 89.90 94.47 96.61
±$\pm 0.45$$\pm 0.03$$\pm 0.25$$\pm 0.02$$\pm 0.00$
RIGA 93.21 90.03 91.23 92.06 93.85
±$\pm 1.12$$\pm 0.91$$\pm 1.21$$\pm 0.58$$\pm 0.09$
LIDC 72.73 72.75 71.83 75.93 82.07
±$\pm 2.70$$\pm 0.66$$\pm 1.50$$\pm 0.53$$\pm 0.92$
In this section, we present a quantitative analysis comparing the performance of our proposed method with various competing approaches across multiple datasets. In Table 5, we report the macro-averaged Dice coefficient, which is calculated as the average of individual class-wise scores. From this aggregate measure, it is seen that our method, enhanced with a confidence regions regularization, consistently outperforms the alternatives. Notably, the standard deviation observed for our approach is generally lower than those of our competitors, indicating the superior stability and reliability of our method.
Detailed tables for each dataset with all six types of Dice coefficients computed separately for each of the classes are provided in Appendix C.1 (Tables 6, 7, 8). It can be seen that our model with a confidence regularizer outperforms the competitors in terms of class-wise Dice coefficients for MNIST dataset. When applied to the LIDC dataset, our approach yields the highest performance with a substantial margin for class 1 (lungs) and second-best result for class 0 (background). Our models show robust performance for classes 0 (background) and 1 (optical disc) on RIGA dataset and achieve the highest score for class 2 (optical cup). In addition, to further illustrate these results, we have included additional visualizations of the final segmentations achieved on each dataset in the Appendix C.1, as depicted in Figures 11, 12, 13 and 14.
Additional experiments, which study the choice and effect of the regularizer, can be found in Appendix C.2.
6 Discussion
Our experimental results demonstrate that the proposed methodologies perform at least as effectively as competing approaches, frequently surpassing them in terms of both the average Dice coefficient (as shown in Table 5) and class-specific Dice scores for segmentation task and in terms of test accuracy for classification task.
A notable observation that appears consistently across all datasets is the superior performance of our model when it uses the proposed confidence regularizer. Moreover, for segmentation, the regularized model shows consistently higher Dice scores than the model without regularization across all classes and datasets.
This consistency of the best performance, regardless of the diversity and complexity of datasets, shows the generalization capability of our model and proposed regularizers, which reassures its potential as a valuable tool for classification and image segmentation tasks of medical data.
7 Conclusion
In this paper, we have introduced a new approach for training classification model with noisy labels and segmentation model under the scenario of multiple noisy annotators. Our key contribution is the introduction of a regularizer which, unlike previous methods, exclusively penalizes the parameters of the base prediction network, thereby incentivizing the model to make more confident predictions. This regulizer allows to improve the results over the competing approaches.
Moreover, we introduce the concept of confident regions for the segmentation problem, identified as regions where all annotators unanimously agree on their predictions, into our loss function. This allows us to push the segmentation model explicitly to predict the class selected by all the annotators. We observed empirically that this consistently helps to improve the results.
Appendix A Proposed architecture for Segmentation
It is worth noting that the ground truth segmentation annotations ${y_{i}}$ are not available to us but assumed to exist. Thus, our goal is to build a model which will approximate the true distribution of class labels $p({y_{i}}\mid x_{i})$ for a given input solely from the dataset of noisy annotations.
A.1 Probabilistic Model for Noisy Observations and Architecture
$p(\{{\tilde{y}^{r}}\}_{r\in{S(x)}}\mid x)$ $=\textstyle{\prod}_{{r\in{S(x)}}}p({\tilde{y}^{r}}\mid x)$ (5)
where ${\tilde{y}^{r}}_{wh}$ is a class label of a pixel in a specific location.
$p({\tilde{y}^{r}}_{wh}\mid x)=\textstyle{\sum}_{l=1}^{L}p({y}_{wh}=l\mid x)\,p({\tilde{y}^{r}}_{wh}\mid{y}_{wh}=l,x).$ (6)
Appendix B Datasets
B.1 Segmentation Datasets

### v3

@@ 5 Experiments
5.1 Classification.
Table 1: Comparison of test accuracy (%) (mean $\pm$ st. dev.) for CIFAR-10. Best results are in bold, second-best underlined. “Conf” stands for confidence regularizer, “Ent” for predictive entropy regularizer.
Noise rate Ours-Conf Ours-Ent Co-tea Co-tea+ JoCoR Trace CDR
@@
Table 2: Comparison of test accuracy (%) (mean $\pm$ st. dev.) for CIFAR-10N. Best results are in bold, second-best underlined. “Conf” stands for confidence regularizer, “Ent” for predictive entropy regularizer.
Noise type Ours-Conf Ours-Ent Co-tea Co-tea+ JoCoR Trace CDR
worst 73.49 $\pm$ 0.61 73.58 $\pm$ 0.96 67.97 $\pm$ 1.17 65.45 $\pm$ 0.25 68.74 $\pm$ 0.21 70.23 $\pm$ 0.89 59.87 $\pm$ 0.58
In this section, we validate our approach to classification problems. For baselines, we will use the following approaches: Co-teaching [13], Co-teaching$+$ [53], JoCoR [46], Robust Early-learning (CDR; [49]) and Annotator Confusion (Trace; [43]). A detailed description of these methods is in the Appendix B.1.
Datasets. In this work, we consider a dataset with synthetically generated annotations along with annotations from human experts. For classification problem, we consider the standard benchmark datasets: MNIST [10], Fashion-MNIST [50], CIFAR-10 [25] where noises are generated synthetically and CIFAR-10N dataset (real-world human annotations noisy data) [47] to demonstrate the effectiveness of our methodology. The noise types, used in the classification experiments, are as follows.
  • The “pairflip noise” involves swapping the labels of two adjacent categories/classes based on a preset ratio [29].
  • The “symmetric noise” retains a portion of the original labels and uniformly reassigns the remainder to all other categories [31].
  • The “worst” noise type (considered exclusively for CIFAR-10N dataset) stands for the highest label noise among human annotations.
Detailed descriptions of datasets and examples of noise transition matrices are provided in Appendix C.1.
[…]
We evaluate the performance of our algorithm in terms of test accuracy for the classifier network. We are particularly interested in the performance of the classifier, as in the evaluation stage this network will be used separately to make predictions. Details on tuning the classification model are in Appendix D.3.
[…]
Table 3: Annotator information for three different styles (MNIST).
Annotators Original Thin Thick
Annotator 1 symmetric 80% asymmetric 40% pairflip 95%
Annotator 3 pairflip 60% pairflip with permutation 40% symmetric 80%
Datasets MR-Net CM-Net PU-NET Ours (no CR) Ours
MNIST 95.81 ±$\pm 0.45$ 93.85 ±$\pm 0.03$ 89.90 ±$\pm 0.25$ 94.47±$\pm 0.02$ 96.61±$\pm 0.00$
RIGA 93.21 ±$\pm 1.12$ 90.03 ±$\pm 0.91$ 91.23 ±$\pm 1.21$ 92.06 ±$\pm 0.58$ 93.85 ±$\pm 0.09$
LIDC 72.73 ±$\pm 2.70$ 72.75 ±$\pm 0.66$ 71.83 ±$\pm 1.50$ 75.93 ±$\pm 0.53$ 82.07 ±$\pm 0.92$
Instance dependent noise and Curated Dataset. In this section, we use our proposed method in an important setup of instance-dependent noise. For this, we have created a dataset that is based on MNIST. Using Morpho-MNIST software [8] we made morphological transformations of images, resulting in different digit styles. We refer to this dataset as Curated MNIST.
Specifically, each image in the Curated MNIST dataset may be either of three different styles. Namely, Original (no transformations are applied), Thick (the digit’s contours become thicker), and Thin (contours become thinner). See Figure 5.1, leftmost column for the demonstration. We have three different annotators, who assign different noisy labels, based on the input image style, making the noise instance-dependent. This is an important experimental setup, that reflects real-world situations. Moreover, the types of noise introduced by annotators are different (see Table 3).
Image Ground truth Ours (m=2, ($\lambda$=0.01) Ours ($\lambda$=0)
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Ground truth (left column) and predicted confusion matrices for different Annotators using different models: our approach with confidence regularizer ($\lambda=0.01$, m=2) (middle column) and without it ($\lambda=0$) (right column) on Curated MNIST.
(a) Testing accuracy
(b) Testing entropy
Figure 3: Accuracy and Entropy for Curated MNIST test dataset computed over seven different seeds. The bold line is the average, shaded region is std. dev.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Visualization of CR with different dilation radius of uncertainty regions (RIGA Optical Cup). Left col. Annotators’ boundaries, Middle col. Initial (d=0), Right col. Dilated (d=3). Here, uncertainty regions have natural colors, while CR is whitened.
In this experiment, we compare (i) the base classifier model, trained with standard cross-entropy loss (without annotators and regularization), (ii) our approach but without regularization, and (iii) our approach with confidence regularizer (see eq. (4.4)). Each annotator’s model has a similar architecture as the classifier model and takes images as an input.
In Figure 5.1, we demonstrate ground truth and learned confusion matrices. Note, that for Thick digit style, noise rates are significant (see Table 3), and despite it is hardly possible to learn the true annotation matrix (e.g. for Annotator 1 for Thick style noise rate is 95%), the resulting performance for the proposed method (see below) is still the best. In Figure 5.1, we plot the testing accuracy and entropy of the base prediction network. Compared to the usage of a classifier model alone or with no regularizer, we see that our approach with the confidence regularizer is both more accurate and confident. More results are in Appendix D.3.2
5.2 Segmentation.
In this section, we evaluate our proposed approach to segmentation problems. As a baseline, we use the following approaches: CM-NET [54], PU-NET [24] and MR-NET [17]. A detailed description of these methods is in the Appendix B.2.
We refer readers to the Appendix D.2 for additional experiments where we adapt the regularizer from [28] to the segmentation problem and compare it to the approaches above.
“Confident regions” (CR) component in the loss. In this work, we also propose to directly apply the segmentation network $\hat{p}_{\theta}(x_{i})$ for predictions for the pixels with low labeling noise. We construct such confident regions (CR) by a special procedure. First, we find such areas of the annotation masks for which all the annotators assign the same label. Next, we use the heuristic that the pixels near the boundary of this region have a high potential for uncertainty. Because of this, we dilate the uncertainty region (complement of the unanimously segmented region) using the convolution operation with zero bias. The complement of this dilated uncertainty region is now a CR, which we will use in our loss function by adding a special term. The process of selecting the CR is illustrated by Figure 4.
Datasets. For the segmentation problem, we use MNIST [10] as a representative of the dataset with artificial noises, where synthetic noisy annotations are created by applying morphological transformations simulations on assumed ground truth [8]. As an example of a real-world dataset we utilize retinal fundus images for glaucoma analysis (RIGA; [1]) and the lung image database consortium (LIDC; [24]) datasets. Descriptions of these datasets are in Appendix C.2.
Performance Evaluation. To evaluate the segmentation model, we use the Dice coefficient on validation splits and report its mean and standard deviation computed over the last 10 epochs. Note the Dice coefficients are calculated for each sample individually, and then the average value is taken. Although the Dice coefficient between the prediction $\hat{p}_{\theta}(x)$ and true label ${y}$ is more important, we also consider Dice coefficient among the noisy annotations ($D(y_{a},y_{a})$), Dice coefficient among predictions of annotations ($D(\tilde{y}_{a},\tilde{y}_{a})$), Dice coefficient between the annotations’ predictions and noisy labels ($D(\tilde{y}_{a},y_{a})$), Dice coefficient between noisy labels and true labels ($D(y_{a},y_{gt})$), and Dice coefficient between the annotations’ predictions and true label ($D(\tilde{y}_{a},y_{gt})$). More details are provided in Appendix D.1. In the case of medical images, to make the evaluation, we compute the true label by majority voting. We emphasize that they are used only in the evaluation stage, not in the training. Along with these baseline models, we also run the model without involving the confidence region in our framework to see the effect of including it.
Segmentation Results. In this section, we compare the performance of our proposed method with various competing approaches across multiple datasets. In Table 4, we report the macro-averaged Dice coefficient, calculated as the average of individual class-wise scores. From this aggregate measure, it is seen that our method, enhanced with a confidence regions regularization, consistently outperforms the alternatives. Notably, the standard deviation observed for our approach is generally lower than those of our competitors, indicating the superior stability and reliability of our method.
Detailed tables for each dataset with all six types of Dice coefficients computed separately for each of the classes are provided in Appendix D.1. It can be seen that our model with a confidence regularizer outperforms the competitors in terms of class-wise Dice coefficients for the MNIST dataset. When applied to the LIDC dataset, our approach yields the highest performance with a substantial margin for class 1 (lungs) and the second-best result for class 0 (background). Our models show robust performance for classes 0 (background) and 1 (optical disc) on the RIGA dataset and achieve the highest score for class 2 (optical cup). We have included additional visualizations of the final segmentations achieved on each dataset in the Appendix D.1. Additional experiments, which study the effect of the regularizer, can be found in Appendix D.2.
6 Discussion
Our experiments demonstrate that the proposed methodology perform at least as competing approaches, frequently surpassing them in terms of both the average Dice coefficient (as shown in Table 4) and class-specific Dice scores for segmentation tasks and in terms of test accuracy for classification task. A notable observation that appears consistently across all datasets is the superior performance of our model when it uses the proposed confidence or entropy regularizers. Moreover, for segmentation, the regularized model shows consistently higher Dice scores than the model without regularization across all classes and datasets. This consistency of the best performance, regardless of the diversity and complexity of datasets, shows the generalization capability of our model and proposed regularizers, which reassures its potential as a valuable tool for classification and image segmentation tasks of medical data.
7 Conclusion
In this paper, we have introduced a new approach for training the classification model with noisy labels and a segmentation model under the scenario of multiple noisy annotators. Our key contribution is the introduction of a class of regularizers that, unlike previous methods, exclusively penalizes the parameters of the base prediction network, thereby incentivizing the model to make more confident predictions. This regularizers allow us to improve the results over the competing approaches.
Moreover, we introduce the concept of confident regions for the segmentation problem, identified as regions where all annotators unanimously agree on their predictions, into our loss function. This allows us to push the segmentation model explicitly to predict the class selected by all the annotators. We observed empirically in both, synthetic and real world noisy data, that this consistently helps to improve the results.
Acknowledgements
The research was partially supported by the RSF grant 20-71-10135. Asma Ahmed Hashmi is supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the German Federal Ministry of Education and Research.
A Proposed architecture for Segmentation
Figure 5: Illustration of the proposed architecture. The input image is fed into two separate networks – Segmentation and Annotation. Then, their predictions are multiplied to produce the prediction of the noisy (annotators) segmentation masks. Finally, the loss function is computed, summing up all the components.
It is worth noting that the ground truth segmentation annotations ${y_{i}}$ are not available to us but assumed to exist. Thus, our goal is to build a model that will approximate the true distribution of class labels $p({y_{i}}\mid x_{i})$ for a given input solely from the dataset of noisy annotations.
The segmentation architecture combines a U-Net [38] based segmentation network with a confusion matrix network to handle noisy labels from multiple annotators. It uses multiple encoder-decoder blocks with skip connections, allowing it to capture both fine and coarse details of the input, while the confusion matrix layers model the noise in the annotations. This is a flexible, highly customizable architecture designed for segmentation tasks, especially in scenarios where annotator label noise is a concern. For comparison with other baselines methods (described in Appendix B.2, we used the same segmentation network.
A.1 Probabilistic Model for Noisy Observations and Architecture
(A.1) $p(\{{\tilde{y}^{r}}\}_{r\in{S(x)}}\mid x)$ $=\textstyle{\prod}_{{r\in{S(x)}}}p({\tilde{y}^{r}}\mid x)$
where ${\tilde{y}^{r}}_{wh}$ is a class label of a pixel in a specific location.
(A.2) $p({\tilde{y}^{r}}_{wh}\mid x)=\textstyle{\sum}_{l=1}^{L}p({y}_{wh}=l\mid x)\,p({\tilde{y}^{r}}_{wh}\mid{y}_{wh}=l,x).$
B Baselines description
B.1 Classification Task Baselines
In this section, we discuss baselines used in comparison in classification problems.
@@ B.1 Classification Task Baselines
B.2 Segmentation Task Baselines
In this section, we discuss baselines used in comparison in segmentation problems.
@@ B.2 Segmentation Task Baselines
C Datasets
C.1 Classification Datasets
MNIST.
The dataset comprises 60,000 samples for training and 10,000 data samples reserved for testing. The number of classes in the dataset is 10.
CIFAR-10.
The CIFAR-10 dataset contains 60,000 color images in 10 classifications, with 6000 images in each class. There are 50,000 training and 10,000 test images. The dataset is divided into five training batches and one test batch, each containing 10,000 images. The test batch is a collection of exactly 1,000 data samples randomly selected from each class. The training batches comprise the remaining images in a random order, however, certain training batches may likely have more images from one class than another.
CIFAR-10N.
CIFAR-10N is prepared by augmenting the training datasets of CIFAR-10 using Amazon Mechanical Turk-collected human-annotated real-world noisy labels. It has the same number of classes and training and test data points as the CIFAR-10 dataset.
FMNIST.
C.2 Segmentation Datasets